OpenStudy (anonymous):

********** metal *********** use a calculator to find each trigonometric ratio. round to the nearest hundredth 2. tan 67= 4 .sin 49= 6 cos 12=

OpenStudy (anonymous):

make sure you in degrees mode not radians. First

OpenStudy (anonymous):

allright

OpenStudy (anonymous):

tan=opp/adj sin=opp/hyp cos=adj/hyp

OpenStudy (anonymous):

yeah i noe

OpenStudy (anonymous):

so if you get tan 67=2.35585 is the same as 2.35585/1 adj=1

OpenStudy (anonymous):

can u answer the other two @timo86m

OpenStudy (anonymous):

what is so hard? it is whatever the calculator gives you over 1? :)

OpenStudy (anonymous):

please just asnwer it @timo86m

OpenStudy (anonymous):

well since it says use calculator use this http://www.wolframalpha.com/input/?i=2.3558523658237528339395866623438831874418383324112790&lk=1&a=ClashPrefs_*Math- just copy past the q

OpenStudy (anonymous):

2. .754/1 3. .978/1